OpenStudy (anonymous):

Represent a right triangular prism in which the height is 100 cm and which base is a right triangle with hypotenuse 30cm in length and an angle of 60 degrees

OpenStudy (mayankdevnani):

height(h) = 100 cm hypotenuse = 30 cm angle = 60 degree

OpenStudy (mayankdevnani):

Area = (1/2)(30sin60)(30cos60) \[\frac{1}{2}\times 30 \sin 60\times 30 \cos 60\]

OpenStudy (mayankdevnani):

Area =194.86 cm^2 \[194.86 cm^2\]

OpenStudy (anonymous):

Is question to find out the area??

OpenStudy (anonymous):

or volume?

OpenStudy (mayankdevnani):

volume =l*b*h \[l \times b \times h\]

OpenStudy (mayankdevnani):

\[l \times b=area\]

OpenStudy (anonymous):

Question is to represent that Triangular Prism I guess..

OpenStudy (mayankdevnani):

volume=100 x 194.86 = 19486 cm^3 \[19486 cm^3\]

OpenStudy (mayankdevnani):

ok @Reyjuanx10

OpenStudy (mayankdevnani):

right @sauravshakya and @waterineyes